OpenStudy (anonymous):

The ratio of the heights of two similar pyramids is 3:2. The volume of the larger pyramid is 135cm^3. What is the volume of the smaller pyramid? A)40cm^3 B)90cm^3 C)27cm^3 D)60cm^3

OpenStudy (anonymous):

volume is \[length^3\] so \[(3/2)^3 = 135/x\] \[27/8 = 135/x\] \[ x = 40cm^3\]